Investors expect Microtech to pay the first dividend of $1.25 at the end of Year 2. The dividend should grow at a rate of 8 percent per year during Years 3, 4, and 5. The stock can be sold for $42 at the end of Year 5. Determine the current market value of the company’s stock if the required rate of return is 12 percent.

Please show work
Answer #1
Price of Stock today is the present value of inflows received during the five years of life:
Therefore, the present value is computed as under:
Year Inflows PVf at 12% Present value
1 0 0.892857 0
2 1.25 0.797194 0.996492
3 1.35 0.71178 0.960903
4 1.458 0.635518 0.926585
5 1.57464 0.567427 0.893493
5 42 0.567427 23.83193
Present value 27.61
Therefore, the price of stock today is $ 27.61
Note: Dividend for Year 3 = Dividend of Year2 i.e. 1.25 + 8% =1.35
Dividend for Year-4 = Dividend for Year-3 + 8% = 1.35+8% = 1.458
Dividend for Year-5 = Dividend for Year-4 + 8% = 1.458+8% = 1.57464